OpenStudy (anonymous):

Two identical, closely wound, circular coils, each having self inductance L, are placed next to each other, so that they are coaxial and almost touching. If they are connected in series, what is the self inductance of the combination? What if they are connected in parallel? Can they be connected so that the total inductance is zero? Explain.

OpenStudy (ghazi):

in parallel equivalent inductance is equal to 1/L=(L1+L2-2M)/L1*L2-M^2 , you have to include mutual inductance and in series L= L1+L2+2M